# Benchmarking Multi-Agent Preference-based Reinforcement Learning for Human-AI Teaming

### v1

@@ A.3 Modified PEBBLE Algorithm for Human Flexibility experiments
Appendix B Experiment Details
B.1 State Space and Action Space
We use the available MuJoCo implementation packaged by OpenAI Gym [62] to instantiate our environments. We use the default state space representation given by the package, and the same observation space is shared by both, the human and the AI agents. The details about the action space are given in Table 1.
For MA Highway, we use the Highway environment packages provided by [11] to create our custom multi-agent environments and the associated tasks. We use the continuous action space consisting of (steering, throttle) for both the agent cars.
B.2 Reward Architecture and Embedding Space
Following the implementation of [15, 13, 14] we implement reward model via a neural network and bound the final output using a $\tanh$ activation function : [-1, 1]. For all the experiments, the reward model has three hidden layers with 256 neurons each followed by the prediction layer (with one neuron). We use the ADAM optimizer for training the SAC actor-critic as well as the reward model. The hyperparameters used for PEBBLE, RUNE & SURF are given in Tables 3, 4, and 5.
The input to the reward model is (state, action) tuple. As used previously [14], we stack the state and action vectors and treat them as a single input for the reward model.
B.3 Human Policy & Preference Oracles
@@
where, $\tilde{R_{h}}$ is the environment reward being used as the human preference reward function and $\tau_{0},\tau_{1}$ are the queried trajectory pairs. Note that we work under the setting that the preference feedback is binary and therefore if the trajectory returns are equal we uniformly pick a preferred trajectory. This does not pose any problems with our chosen benchmark domains as the underlying reward is dense and shaped [63, 64].
@@
B.4 Implementation, Code and Compute
We use the publicly available implementation of [15, 13, 14] for the implementation of SAC, PEBBLE, RUNE & SURF algorithms. Our implementation of extending these PbRL algorithms to the multi-agent setting can be found in the code given in the supplementary. All the experiments were run on an Intel(R) Xeon(R) Gold 6258R CPU @ 2.70GHz, with Quadro RTX 8000 GPU.
Appendix C Hyperparameters
@@
50 (MA MuJoCo) Optimizer Adam [65]

### v2

@@ Appendix C Experiment Details
We use the available MuJoCo implementation packaged by OpenAI Gym (Brockman et al. 2016) to instantiate our environments. We use the default state space representation given by the package, and the same observation space is shared by both, the human and the AI agents. The details about the action space are given in Table 4.
For MA Highway, we use the Highway environment packages provided by (Leurent 2018) to create our custom multi-agent environments and the associated tasks. We use the continuous action space consisting of (steering, throttle) for both the agent cars.
Reward Architecture and Embedding Space
Following the implementation of (Lee, Smith, and Abbeel 2021; Liang et al. 2022; Park et al. 2022) we implement reward model via a neural network and bound the final output using a $\tanh$ activation function : [-1, 1]. For all the experiments, the reward model has three hidden layers with 256 neurons each followed by the prediction layer (with one neuron). We use the ADAM optimizer for training the SAC actor-critic as well as the reward model. The hyperparameters used for PEBBLE, RUNE & SURF are given in Tables 6, 7, and 8.
The input to the reward model is (state, action) tuple. As used previously (Park et al. 2022), we stack the state and action vectors and treat them as a single input for the reward model.
Human Policy & Preference Oracles
@@
where, $\tilde{R_{h}}$ is the environment reward being used as the human preference reward function and $\tau_{0},\tau_{1}$ are the queried trajectory pairs. Note that we work under the setting that the preference feedback is binary and therefore if the trajectory returns are equal we uniformly pick a preferred trajectory. This does not pose any problems with our chosen benchmark domains as the underlying reward is dense and shaped (Devlin and Kudenko 2012; Ng, Harada, and Russell 1999).
@@
Implementation, Code and Compute
We use the publicly available implementation of (Lee, Smith, and Abbeel 2021; Liang et al. 2022; Park et al. 2022) for the implementation of SAC, PEBBLE, RUNE & SURF algorithms. Our implementation of extending these PbRL algorithms to the multi-agent setting can be found in the code given in the supplementary. All the experiments were run on an Intel(R) Xeon(R) Gold 6258R CPU @ 2.70GHz, with Quadro RTX 8000 GPU.
Appendix D Hyperparameters
@@
50 (MA MuJoCo) Optimizer Adam
@@
Appendix E Additional Results
The learning curves on MA Highway-Left and MA Highway-Middle have been shown in Figure 19.
Figure 19: Learning curves on MA Highway-Left (row 1) and MA Highway-Middle (row 2): comparing (from L to R) (a) Human Flexibility on multiple $\pi_{H}$, with Specified Orchestration case that assumes a single $\pi_{H}$ and complete access to it, (b) agent’s 0% access to $\pi_{H}$, and (c) partial access to $\pi_{H}$; as measured on the episodic returns. The solid lines and shaded regions represent the mean and standard deviation, respectively, across three runs.
Appendix F Societal Impact